OpenStudy (anonymous):

i will medal and fan for help with question David is rowing a boat upstream. The river is flowing at a speed of 2 miles per hour. David starts rowing at a speed of 6 miles per hour, but as he gets tired his speed decreases (at a rate of 1 mile per hour, every hour). Which equation represents the speed of the boat for x hours spent rowing?

OpenStudy (anonymous):

y = 6 − 2x y = 4 − x y = x + 4 y = 6 − x y = x− 4

OpenStudy (anonymous):

y=6-2x y=4-x y=x+4 y=6-x y=x-4

OpenStudy (anonymous):

is their more than 1 answer choice?

OpenStudy (anonymous):

those 5 are the answer choices

OpenStudy (anonymous):

but theres only 1 answer out of all of them correct?

OpenStudy (anonymous):

yea i know i need to know which one of those 5 are my answer

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

DECREASES at a rate of 1 mile per hour This would be subtracting 'x'. He's rowing upstream, the speed of the river is 2MPH, and he's rowing 6 MPH. So what's 6 - 2?

OpenStudy (anonymous):

4

OpenStudy (igreen):

Yes, so we have 4 - x. So our answer is: y = 4 - x

OpenStudy (anonymous):

y=x-4

OpenStudy (igreen):

No, that's close..we subtract 'x' though..not '4'..

OpenStudy (anonymous):

thank you can you help me with another one

OpenStudy (anonymous):

What is the slope and y-intercept of the equation represented by the following graph? slope = 2, y-intercept =-1/2 slope = -2, y-intercept =-1/2 slope =-1/2, y-intercept = 2 slope =1/2 , y-intercept = 2 slope = -1, y-intercept = 2

OpenStudy (igreen):

Sure, just close this one and oepn a new one.